OpenStudy (anonymous):

Jill sold half of her comic books and then bought sixteen more. she now has 36. With how many did she begin?

jimthompson5910 (jim_thompson5910):

Let x = number she had at the beginning

jimthompson5910 (jim_thompson5910):

"Jill sold half of her comic books", so she has x - x/2 = 2x/2 - x/2 = x/2 comics left

jimthompson5910 (jim_thompson5910):

"...and then bought sixteen more." means she now has x/2 + 16 comics

jimthompson5910 (jim_thompson5910):

Since we know she now has 36 books, we know that x/2 + 16 = 36

jimthompson5910 (jim_thompson5910):

Solve that for x to get your answer

OpenStudy (anonymous):

how would i start it off

jimthompson5910 (jim_thompson5910):

subtract 16 from both sides

OpenStudy (anonymous):

then divide both by 2?

jimthompson5910 (jim_thompson5910):

then multiply both sides by 2

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

40

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome